OpenStudy (anonymous):

Please help me Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form. –2x – y = –7 2x – y = –7 2x – y = 11 –2x + y = –11

OpenStudy (nubeer):

u need to know the equation first y-y1 = m (x-x1) m = gradient, x1 = x-coordinate, y1 = y-coordinate just plug in the values u will get the equation

OpenStudy (anonymous):

@MarieSoLoved

OpenStudy (anonymous):

i got 2x-y=-7

OpenStudy (anonymous):

same^

OpenStudy (anonymous):

Need anymore help just find me